# Bi-Lipschitz embedding metric triangles in the plane

### v1

@@ 5 Metric polygons
6 Appendix
In this appendix, we provide the proofs of the various inequalities used in Section 3. These are mostly straightforward calculus optimization problems, but we include the details for completeness. We recall the following functions defined in Section 3:
@@ 6 Appendix
$\frac{\partial M_{1}}{\partial e}(b,e)=-\frac{2+2b+2b^{2}+e-be}{e^{3}}.$
@@
Observe that $\frac{\partial M_{2}}{\partial b}(b,e)>0$ if $12b<e-$ 1 and $\frac{\partial M_{2}}{\partial b}(b,e)<0$ if $12b>e-$ 1. This implies that $M_{2}(b,e)$ is greatest if $2b+1=e$, or if $b=0$.
@@
which is negative. Thus $M_{2}(b,2b+1)$ is maximized when $b=0$, which gives a value of $4$.
@@
which is negative if $e>0$ or $e<-2$ and positive if $-2<e<0$. Thus $M_{2}(0,e)$ has a local maximum at $e=0$. We also check that $\lim_{e\to-\infty}M_{2}(0,e)=1$, and so $e=0$ is a global maximum. This completes the proof.
$M_{2}(-e,e)=\frac{(2-e)^{2}}{\frac{1}{4}(1+3e)^{2}+\frac{3}{4}(1-e)^{2}}.$
$\frac{2(e-2)(1+6e)}{(1+3e^{2})^{2}}.$
@@
The function $\widetilde{M}_{2}(b,e)$ is maximized on the set $\{(b,e):b\geq 0,-b\leq e\leq 0\}$ by taking $b=-e=1/6$, which gives a value of $13/3$.
@@
$\frac{\partial M_{2}}{\partial b}(b,e)=-\frac{(1+2b-e)(e-2)^{2}}{(\frac{1}{4}\left(1-b+2e\right)^{2}+\frac{3}{4}(1+b)^{2})^{2}}.$
This is strictly negative over the given region. It follows that $M_{2}(b,e)$ is maximized if $b=-e$. Now
$\frac{\partial}{\partial e}M_{2}(-e,e)=\frac{2(e-2)(1+6e)}{(1+3e^{2})^{2}}.$
This is positive if $e<-1/6$ and negative if $e>-1/6$. Thus there is a maximum at $e=-1/6$. We evaluate $M_{2}(1/6,-1/6)=13/3$. ∎
@@
$\frac{\partial M_{3}}{\partial s}(b,s,t)=-\frac{(2+2b-s-t)(6+7b+3b^{2}+s-3t)}{2(\frac{1}{4}(1-b+2s-2t)^{2}+\frac{3}{4}(1+b)^{2})^{2}},$
which is negative for all $10<b,s,t<$ 1. It follows that $M_{3}(b,s,t)$ is maximized if $s=0$. Next, we evaluate
$\frac{\partial}{\partial t}M_{3}(b,0,t)=-\frac{(2+2b-t)(2+3b+3b^{2}-t)}{2(\frac{1}{4}(1-b-2t)^{2}+\frac{3}{4}(1+b)^{2})^{2}},$
which is also negative for all $10<b,t<$ 1. Thus $M_{3}(b,0,t)$ is maximized if $t=0$. Finally, we evaluate
$\frac{\partial M_{3}}{\partial b}(b,0,0)=\frac{(3-b)(1+b)}{(\frac{1}{4}(1-b)^{2}+\frac{3}{4}(1+b)^{2})^{2}},$
which is positive if $10<b<$ 1. We conclude that $M_{3}(b,s,t)$ is maximized when $b=1$, $s=t=0$. Evaluate $M_{3}(1,0,0)=16/3$. ∎
@@
Let $u=b-s$, so the previous expression becomes $h(u)=\frac{3}{4}(2+u)^{2}+\frac{1}{4}(2-u)^{2}$. Then $h^{\prime}(u)=2u+2$. So $M_{4}$ is maximized when $b-s$ is greatest, that is, $b=1$ and $s=0$. Then $t=1$. We evaluate $M_{4}(1,0,1)=7$. ∎
@@
which is positive when $12b>s-t-$ 1. If $12b>s-t-$ 1, then $\frac{1}{(s+t)^{2}}M_{4}(b,s,t)$ increases as $b$ increases. If $12b<s-t-$ 1, then $\frac{1}{(s+t)^{2}}M_{4}(b,s,t)$ increases as $b$ deceases. From this we conclude that $\frac{1}{(s+t)^{2}}M_{4}(b,s,t)$ is greatest when either $b=s+t$ or $b=0$. When analyzing the latter case, note that $t+1<s$. Then
@@
Let $u=(\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}})$. We compute
$\frac{\partial}{\partial u}\frac{M_{4}(0,s,t)}{(s+t)^{2}}=\frac{s-t-1}{\sqrt{2}(s+t)^{2}},$
@@
Let $u=(0,\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$. The directional derivative of $\frac{1}{b^{2}}M_{4}(b,s,t)$ in the direction of $u$ is $3(1+s+t)/(\sqrt{2}b^{2})$, which is positive. We conclude that $\frac{1}{b^{2}}M_{4}(b,s,t)$ is greatest when $s+t=b$. This reduces to 6.7, and so we obtain the same conclusion. ∎
@@
Let $M_{5}^{a}(b,s,t)$ denote the function defined like $M_{5}$ but replacing “$\min\{2,2b\}$” with “$2$”, and similarly for $M_{5}^{b}$. It suffices to consider the functions $M_{5}^{a}$ and $M_{5}^{b}$ separately.
@@
This is positive for $b\geq 0$, $10<s<$ 1, so we conclude that $s=1$.
@@
This is positive if $2s<-1+b-t$ and negative if $2s>-1+b-t$. Thus any maximum of $M_{6}$ occurs when $2s=-1+b-t$, if $-1+b-t\geq 0$, or $-1+b-t<0$ and $s=0$. In the first case, we evaluate $M_{6}(b,(-1+b-t)/2,t)=16/3$.
@@
If $-1+b-t<0$, then this derivative is positive. We conclude that any maximum for $M_{6}(b,0,t)$ occurs when $-1+b-t=0$, which returns us to the previous case. ∎
@@
which is positive for $10<b,s,t<$ 1. This implies the function is greatest when $b=1$. We now compute
@@
If $s-t<1-b$, then $D_{u}M_{5}$ is positive; if $s-t>1-b$, then $D_{u}M_{5}$ is negative. Thus along a given curve in the $u$-direction, $M_{8}(b,s,t)$ is maximized when $s=1-b+t$. Note that this point may lie outside the target region $\{(b,s,t):0\leq b,s,t\}$, in which case the maximum necessarily occurs when $t=0$. We examine each case separately. First, we take $s=1-b+t$. Then
@@
The function $M_{9}(b,s,t)$ is maximized on the set $\{(b,s,t):0\leq b,s,t\}$ by taking $t=b+s-1$, which gives a value of $16/3$.
@@
This is positive if $t+1<b+s$ and negative if $t+1>b+s$, so we conclude that $M_{9}(b,s,t)$ is largest when $t=b+s-1$. Now $M_{9}(b,s,b+s-1)$ evaluates to $16/3$. ∎

### v2

@@ 5. Metric polygons
6. Appendix
In this appendix, we provide the proofs of the various inequalities used in Section˜3. These are mostly straightforward calculus optimization problems, but we include the details for completeness. We recall the following functions defined in Section˜3:
@@ 6. Appendix
$\frac{\partial}{\partial e}\frac{M_{1}(b,e)}{e^{2}}=-\frac{2+2b+2b^{2}+e-be}{e^{3}}.$
@@
Observe that $\frac{\partial M_{2}}{\partial b}(b,e)>0$ if 2b<e12b<e-1 and $\frac{\partial M_{2}}{\partial b}(b,e)<0$ if 2b>e12b>e-1. This implies that $M_{2}(b,e)$ is greatest if $2b+1=e$, or if $b=0$.
@@
which is negative. Thus $M_{2}(b,2b+1)$ is maximized when $b=0$, which gives a value of $M(0,1)=4/3$.
@@
which is negative if $e>0$ or e<2e<-2 and positive if 2<e<0-2<e<0. Thus $M_{2}(0,e)$ has a local maximum at $e=0$, which gives the value $M_{2}(0,0)=4$. We also check that $\lim_{e\to-\infty}M_{2}(0,e)=1$, and so $e=0$ is a global maximum. This completes the proof. ∎
@@
The function $\widetilde{M}_{2}(b,e)$ is maximized on the set $\{(b,e):b\geq 0,-b\leq e\leq 0\}$ by taking b=e=1/6b=-e=1/6, which gives a value of $13/3$.
@@
$\frac{\partial\widetilde{M}_{2}}{\partial b}(b,e)=-\frac{(1+2b-e)(e-2)^{2}}{(\frac{1}{4}\left(1-b+2e\right)^{2}+\frac{3}{4}(1+b)^{2})^{2}}.$
This is strictly negative over the given region. It follows that $\widetilde{M}_{2}(b,e)$ is maximized if b=eb=-e. Now
$\frac{\partial}{\partial e}\widetilde{M}_{2}(-e,e)=\frac{2(e-2)(1+6e)}{(1+3e^{2})^{2}}.$
This is positive if e<1/6e<-1/6 and negative if e>1/6e>-1/6. Thus there is a maximum at e=1/6e=-1/6. We evaluate $\widetilde{M}_{2}(1/6,-1/6)=13/3$. ∎
@@
$\frac{\partial M_{3}}{\partial s}(b,s,t)=-\frac{(2+2b-s-t)(s(5-4t)+b(2+3s-t)+4(t-\frac{7}{8})^{2}+\frac{15}{16})}{(\frac{1}{4}(1-b+2s-2t)^{2}+\frac{3}{4}(1+b)^{2})^{2}},$
which is negative for all $0<b,s,t<1$. It follows that $M_{3}(b,s,t)$ is maximized if $s=0$. Next, we evaluate
$\frac{\partial}{\partial t}M_{3}(b,0,t)=-\frac{(2+2b-t)(4b^{2}+b(5t+2)+3t)}{2(\frac{1}{4}(1-b-2t)^{2}+\frac{3}{4}(1+b)^{2})^{2}},$
which is also negative for all $0<b,t<1$. Thus $M_{3}(b,0,t)$ is maximized if $t=0$. Finally, we evaluate
$\frac{\partial M_{3}}{\partial b}(b,0,0)=\frac{4(1-b^{2})}{(\frac{1}{4}(1-b)^{2}+\frac{3}{4}(1+b)^{2})^{2}},$
which is positive if $0<b<1$. We conclude that $M_{3}(b,s,t)$ is maximized when $b=1$, $s=t=0$. Evaluate $M_{3}(1,0,0)=16/3$. ∎
@@
Let u=bsu=b-s, so the previous expression becomes $h(u)=\frac{3}{4}(2+u)^{2}+\frac{1}{4}(2-u)^{2}$. Then $h^{\prime}(u)=2u+2$. So $M_{4}$ is maximized when bsb-s is greatest, that is, $b=1$ and $s=0$. Then $t=1$. We evaluate $M_{4}(1,0,1)=7$. ∎
@@
which is positive when 2b>st12b>s-t-1. If 2b>st12b>s-t-1, then $\frac{1}{(s+t)^{2}}M_{4}(b,s,t)$ increases as $b$ increases. If 2b<st12b<s-t-1, then $\frac{1}{(s+t)^{2}}M_{4}(b,s,t)$ increases as $b$ decreases. From this we conclude that $\frac{1}{(s+t)^{2}}M_{4}(b,s,t)$ is greatest when either $b=s+t$ or $b=0$. When analyzing the latter case, note that $t+1<s$. Then
@@
Let $u=(0,\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}})$. We compute
$D_{u}\frac{M_{4}(0,s,t)}{(s+t)^{2}}=\frac{s-t-1}{\sqrt{2}(s+t)^{2}},$
@@
Let $u=(0,\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$. The directional derivative of $\frac{1}{b^{2}}M_{4}(b,s,t)$ in the direction of $u$ is $3(1+s+t)/(\sqrt{2}b^{2})$, which is positive. We conclude that $\frac{1}{b^{2}}M_{4}(b,s,t)$ is greatest when $s+t=b$. This reduces to ˜6.7, and so we obtain the same conclusion. ∎
@@
Let $M_{5}^{a}(b,s,t)$ denote the function defined like $M_{5}$ but replacing $\min\{2,2b\}$” with “22”, and similarly for $M_{5}^{b}$. It suffices to consider the functions $M_{5}^{a}$ and $M_{5}^{b}$ separately.
@@
This is positive for $b\geq 0$, $0<s<1$, so we conclude that $s=1$.
@@
This is positive if 2s<1+bt2s<-1+b-t and negative if 2s>1+bt2s>-1+b-t. Thus any maximum of $M_{6}$ occurs when 2s=1+bt2s=-1+b-t, if $1+b-t\geq 0$, or 1+bt<0-1+b-t<0 and $s=0$. In the first case, we evaluate $M_{6}(b,(-1+b-t)/2,t)=16/3$.
@@
If 1+bt<0-1+b-t<0, then this derivative is positive. We conclude that any maximum for $M_{6}(b,0,t)$ occurs when 1+bt=0-1+b-t=0, which returns us to the previous case. ∎
@@
which is positive for $0<b,s,t<1$. This implies the function is greatest when $b=1$. We now compute
@@
If st<1bs-t<1-b, then $D_{u}M_{5}$ is positive; if st>1bs-t>1-b, then $D_{u}M_{5}$ is negative. Thus along a given curve in the $u$-direction, $M_{8}(b,s,t)$ is maximized when s=1b+ts=1-b+t. Note that this point may lie outside the target region $\{(b,s,t):0\leq b,s,t\}$, in which case the maximum necessarily occurs when $t=0$. We examine each case separately. First, we take s=1b+ts=1-b+t. Then
@@
The function $M_{9}(b,s,t)$ is maximized on the set $\{(b,s,t):0\leq b,s,t\}$ by taking t=b+s1t=b+s-1, which gives a value of $16/3$.
@@
This is positive if $t+1<b+s$ and negative if $t+1>b+s$, so we conclude that $M_{9}(b,s,t)$ is largest when t=b+s1t=b+s-1. Now $M_{9}(b,s,b+s-1)$ evaluates to $16/3$. ∎
Acknowledgments: We thank Guy C. David and Dustin Mixon for their feedback and corrections on a draft of this paper. We also thank the anonymous referees for their comments and corrections.
Funding information: M.R. was partially supported by the National Science Foundation under grant DMS-2413156.